OpenStudy (anonymous):

Simplify the difference quotient [f(x+h) - f(x)]/h, where f(x)=6x+5

OpenStudy (anonymous):

this is the same as asking for the slope it is the change in \(y\) over the change in \(x\) should be clear before you start since you know the slope of \(y=6x+5\)

OpenStudy (anonymous):

of course if you like you can write \[\frac{f(x+h)f(x)}{h}=\frac{6(x+h)+5-(6x+5)}{h}\] so some algebra and you will get the answer